An ice cube at 0 oC is dropped on a ground and it melts to water at 0 oC. If all the kinetic energy of ice went into melting it, from what height did it fall
1
Expert's answer
2019-06-04T12:12:53-0400

Beginning potential energy of the ice cube: "E_p=m*g*h"

Where,

"m -" weight of ice cube

"h -" height of drop

"g -" gravitational acceleration ("g=9.8 m\/s^2" for SI)


At the monet of the falling end kinetic energy: "E_k=m*v^2\/2"

Where,

"v-" speed in the moment of contact with groung

Using the energy conversation law:"E_k=E_p"

Using the heat of fusion formula: "Q=L*m"

Where,

"Q-" nesessary heat for melting

"L-" heat of fusion of ice ("L=334*10^3 J\/kg" for SI).

According to the task: "E_k=Q" or "L*m=m*g*h" "\implies" "h=L\/g"

Using numbers, "h=34.08*10^3 m"

Answer: "34.08*10^3 m"